Two samples consisting of 21 and 9 observations have variances given by s 1 2 =16 and s 2 2

=8 respectively. Test the hypothesis that the first population variance is greater than the

second at a (a) 0.05, (b) 0.01 level of significance.


1
Expert's answer
2022-03-03T10:40:05-0500

The provided sample variances are "s_1^2=16" and "s_2^2=8"  and the sample sizes are given by "n_1=21" and "n_2=9."

The following null and alternative hypotheses need to be tested:

"H_0:\sigma_1^2=\sigma_2^2"

"H_1:\sigma_1^2>\sigma_2^2"

This corresponds to a right-tailed test, for which a F-test for two population variances needs to be used.

(a) Based on the information provided, the significance level is "\alpha=0.05," and the rejection region for this right-tailed test is "R=\{F:F>F_U=3.15\}."

The F-statistic is computed as follows:



"F={s_1^2\over s_2^2}={16\over 8}=2"

Since from the sample information we get that "F=2\leq3.15=F_U," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population variance "\sigma_1^2" is greater than the population variance "\sigma_2^2," at the "\alpha=0.05" significance level.


(b) The significance level is "\alpha=0.01," and the rejection region for this right-tailed test is "R=\{F:F>F_U=5.36\}."

The F-statistic is computed as follows:


"F={s_1^2\over s_2^2}={16\over 8}=2"

Since from the sample information we get that "F=2.25\leq5.36=F_U," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population variance "\sigma_1^2" is greater than the population variance "\sigma_2^2," at the "\alpha=0.01" significance level.